GROUPS AS OF FRIDAY, FEB. 19

A ALPHA:  Brown, Cobb, Murray
B BETA:  Chaney, King
Γ GAMMA:  Clough, Ward
Δ DELTA:  Combest, Generes
E EPSILON:  Denniston, Guerrero
Z ZETA:  HaleAl, Webster
H ETA:  Elmquist, Ibarra
Θ THETA:  Glenn, Jones, Peterson
I IOTA:  Frerich, Klaus, Runyan
K KAPPA:  Lam, Modjdehi, NguyenQ
Λ LAMBDA:  Davis, Linton
M MU:  HaleAn, Kiser, Thomson
N NU:  Beason, Ninh, Schleicher
Ξ XI:  Aguirre, Mendoza, NguyenN
O OMICRON: Potts, Wood
Π PI:  Cervenka, NguyenM, O’Brien
P RHO:  Kouri, Noyes, Spence
Σ SIGMA:
T TAU:
Υ UPSILON:
Φ PHI:
X CHI:
Ψ PSI:
Ω OMEGA: